Solve the Non-exact Equation



v(u2+v2)du - u(u2+2v2)dv

1
Expert's answer
2022-02-02T09:23:43-0500

Solution:

Given, "v(u^2+v^2)du - u(u^2+2v^2)dv=0"

Replace u with x and v with y.

"y(x^2+y^2)dx - x(x^2+2y^2)dy=0
\\ \Rightarrow \dfrac{dy}{dx}=\dfrac{y(x^2+y^2)}{x(x^2+2y^2)}\ ...(i)"

Put "y=vx\ ...(ii)"

"\Rightarrow \dfrac{dy}{dx}=v+x\dfrac{dv}{dx}\ ...(iii)"

From (i), (ii) and (iii).

"v+ \dfrac{dv}{dx}=\dfrac{vx(x^2+v^2x^2)}{x(x^2+2v^2x^2)}
\\ \Rightarrow \dfrac{dv}{dx}=\dfrac{v(1+v^2)}{(1+2v^2)}-v
\\ \Rightarrow \dfrac{dv}{dx}=\dfrac{v+v^3-v-2v^3}{(1+2v^2)}
\\ \Rightarrow \dfrac{dv}{dx}=\dfrac{-v^3}{(1+2v^2)}
\\ \Rightarrow \dfrac{(1+2v^2)dv}{v^3}=-dx
\\ \Rightarrow (\dfrac{1}{v^3}+\dfrac{2}{v})dv=-dx"

On integrating both sides,

"-\dfrac{1}{2v^2}+2\log v=-x+C
\\ \Rightarrow-\dfrac{1}{2(\frac yx)^2}+2\log \frac yx=-x+C
\\ \Rightarrow-\dfrac{x^2}{2y^2}+2\log \dfrac yx=-x+C"

Now, replace u with x and v with y back

"-\dfrac{u^2}{2v^2}+2\log \dfrac vu=-u+C"